OpenStudy (anonymous):

11. Which of the following describes all the solutions to the inequality 5𝑥 + 7𝑦 ≥ 22 when 𝑦 = −4? A. 𝑥 ≤ 10 B. 𝑥 ≤ −10 C. 𝑥 ≥ 10 D. 𝑥 ≥ −10

Directrix (directrix):

5𝑥 + 7𝑦 ≥ 22 --> Replace y by -4 and solve the inequality. @mdfrmhvn

OpenStudy (anonymous):

-28

Directrix (directrix):

Not correct.

OpenStudy (anonymous):

woud you subtract -28 from both sides

Directrix (directrix):

5𝑥 + 7𝑦 ≥ 22 5x + 7*(-4) >= 22 5x -28 >= 22 5x >= 22 + 28 Solve for x.

OpenStudy (anonymous):

o 50 / 5 =10

Directrix (directrix):

x does not equal 10.

Directrix (directrix):

5x >= 50 x >= ?

OpenStudy (anonymous):

x>=10 im not understanding how x isnt 10 when my answer choices all have 10

Directrix (directrix):

You have equalities and inequalites mixed up. x is 10 but x is also 10 AND all numbers greater than 10.

Directrix (directrix):

Hence, x>= 10

OpenStudy (anonymous):

well now i understand

Directrix (directrix):

That is good news.

OpenStudy (anonymous):

really thank you i appriciate it i have a hard time learning

Directrix (directrix):

Keep working. Don't ever give up.

OpenStudy (anonymous):

:) thank you